OpenStudy (anonymous):

Determine which ordered pair is a solution of y=4x+2. a. (4, 14) b. (–1, –2) c. (10, 2) d. (–2, 10)

OpenStudy (anonymous):

just sub in your answer choices and see which one is true... y = 4x + 2 (4,14)..x = 4 and y = 14 14 = 4(4) + 2 14 = 16 + 2 14 = 18 (not this one because it is not true) y = 4x + 2 (-1,-2) -2 = 4(-1) + 2 -2 = -4 + 2 -2 = -2 (found it....this one is true) answer is (-1,-2)

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

anytime :)